# Jaime Hernández-Lasheras

### v1

@@ Abstract
The impact of the expansion of a high-frequency radar (HFR) system in the Ibiza Channel (Western Mediterranean Sea) is evaluated through an Observing System Simulation Experiment (OSSE). The installation of two new antennas in the Iberian Peninsula would complement the existing ones in the islands of Ibiza and Formentera, providing surface currents observations of the full channel. Two different configuration of the same model, validated to give realistic simulations, are used: i) a Nature Run (NR) which is considered as the real ocean state and that is used to generate pseudo-observations, and ii) a Control Run (CR) in which we will assimilate the pseudo-observations. The OSSE is first validated by comparison against a previous Observing System Experiment (OSE). The effect of the new antennas for forecasting surface currents is evaluated in two different periods. The effect of the new antennas is relatively small when the NR and CR depict a similar circulation. However, in situations where both models present higher differences, the error reduction with respect to the use of only the actual system can be of up to 19%. The effects on the transport in the area are also analyzed from a Lagrangian perspective, showing that DA can help to better represent the Lagrangian Coherent Structures present in the NR and constrain the ocean dynamics.
@@ 1 Introduction
Numerical models provide a complete view of the three dimensional structure of the ocean. However, they are inevitably affected by errors from parametrization of non resolved physical processes, discretization issues, or the lack of accurate forcing. To improve reliability, numerical models for operational purposes should be fed with observations through data assimilation. Observing System Experiments (OSEs) assimilating data are numerical experiments that can be designed to evaluate the capability of specific observing systems and eventually correct model forecasts on simulations. Similarly, the potential impact of observing system has to be evaluated to help design these systems. Observing System Simulation Experiments (OSSEs) can be performed to help to optimally design an OOS or a future campaign (Kourafalou et al., 2015).
OSSEs were first developed for the atmospheric science community, and over time, specific design criteria have been developed to ensure the realism of the assessments performed, as defined in Atlas (1997). In ocean studies multiple OSSEs had been done, however, most of them did not use a full-fledged DA system approach for the evaluation. Generally, Kalman filters, empirical orthogonal functions (EOF) based or different interpolation methods were used to map the observations and reconstruct the ocean state (Guinehut et al., 2004; Ballabrera-Poy et al., 2007; Sakov and Oke, 2008). Following the procedure established for atmospheric studies (Hoffman and Atlas, 2016), Halliwell et al. (2014) applied them for the first time in the ocean, and in the last years, several studies have been done following the criteria exposed in that work, as we will do here. For instance, Gasparin et al. (2019) performed an evaluation of the influence of the future deep Argo float network, and Benkiran et al. (2021) assessed the impact of the assimilation of data from the future SWOT satellite mission in a global-high-resolution model.
@@
Normally, the validation of these Lagrangian techniques is limited. LCS computed from model simulations can be compared to those calculated from geostrophic currents derived from altimetry products (Hernández-Carrasco and Orfila, 2018) which suffer limitations when approaching the coast (Vignudelli et al., 2019; Pascual et al., 2013). Also from HFR measured currents (Hernández-Carrasco et al., 2018), that are limited to cover small coastal areas. The validation of LCS can be performed by comparison with active tracers, as chlorophyll filaments (Lehahn et al., 2007; Hernández-Carrasco et al., 2018), which by their nature can not depict the full structures present in the ocean; SST fronts (d’Ovidio et al., 2004), which are inferred from satellite products that can be affected by clouds; or fish stock concentrations (Díaz-Barroso et al., 2022), tracked seabirds or marine predators (Kai et al., 2009), that are difficult to monitor. Here we will take profit of the full ocean state knowledge supposed in the OSSE approach. The use of a NR model for the validation of the LCS computed from the model simulations implies a step forward to address the question of how data assimilation can help models to correct the circulation, especially in coastal areas.
This study is a continuation of the work presented in Hernandez-Lasheras et al. (2021), where a series of OSEs were performed to evaluate the impact of HFR DA on the correction of surface currents in the Ibiza Channel (IC). In that work, it was shown that using HFR observations together with satellite observations (altimetry and sea surface temperature) and Argo temperature and salinity profiles increased the model’s capability to forecast surface currents. In this work we will use the same set-up, using the nudging initialization method described in the mentioned work, as it is the configuration employed in the WMOP, the Western Mediterranean OPerational forecasting system.
@@ 2.1 Study Area and HFR system
SOCIB operates since 2012 two CODAR HFR antennas in the islands of Ibiza and Formentera, measuring surface currents in an area up to 80 km far off the coast (Tintoré et al., 2012). Here we evaluate the potential impact that a couple of new antennas eventually installed in the Iberian peninsula, in the eastern part of Cape la Nao (Fig. 1), would have in the WMOP system. By installing these new antennas, the IC HFR system will expand its coverage area covering the entire channel. The area, shown in Fig. 1 (red dashed line), is considered as the most likely coverage in terms of total velocities (u-v) that a couple of antennas in the peninsula will provide together with the actual system.
Figure 1: WMOP domain and bathymetry in the Western Mediterranean. HFR actual coverage area (black dashed) and Future one (red dashed) from which observations have been simulated. The area selected for validation is also shown (blue).
@@ 2.2 OSSE set-up: Simulations
The OSSE perspective requires two model simulations. Here we used a fraternal twin OSSE system approach (Halliwell et al., 2014), in which two different configurations of the same model are employed: i) a Nature Run (NR), which is considered to represent the true ocean and that will be used for validation and to extract the pseudo-observations, and (ii) a Control Run (CR) model in which we will evaluate the impact of assimilating different observing sources. We use the WMOP (Juza et al., 2016) model, which is a regional configuration of ROMS (Shchepetkin and McWilliams, 2005) for the Western Mediterranean Sea. It spans from Gibraltar Strait in the west to Corsica and Sardinia straits in the east, with a horizontal resolution around 2 km and 32 vertical terrain-following sigma levels (resulting in a vertical resolution between 1 and 2m at the surface). Both the NR and the CR model, which will be used to assimilate, have the same configuration, parametrization and atmospheric forcing, with the only difference in the initial state and the boundary conditions used in each of them. The CR is a free-run hindcast simulation developed and evaluated in Mourre et al. (2018); Aguiar et al. (2019). We will use this model configuration to assimilate data into, as it is the same one used in the reference OSE we will use for validation (Hernandez-Lasheras et al., 2021). It uses the Copernicus Marine Forecasting System for the Mediterranean Sea (CMEMS MED-MFC), with a 1/16 horizontal resolution (Simoncelli et al., 2017), as initial and lateral boundary conditions. NR, in contrast, uses the Mercator Glorys reanalysis global product, with a 1/12 horizontal resolution (CMEMS GLOBAL_REANALYSIS_PHY_001_030) and has also been validated to give realistic simulations, comparing against observational data from satellites and Argo buoys. The atmospheric forcing, common for both simulations, is provided every 3 hours at $1/20^{\circ}$ resolution by the Spanish Meteorological Agency (AEMET) through the HIRLAM model (Undén et al., 2002) and the bathymetry is derived from a 1’ database (Smith and Sandwell, 1997).
@@
To further explore the capabilities of the new antenna under different possible circumstances, we selected another period in which the dynamics in the area from the two simulations presented more differences. For this, we chose August 2014 to repeat our simulations. In the CR simulation, the dynamics during August are similar to the following September-October period, with northward currents in the east side of the channel and southward in the west, as can be seen in Fig. 2. On the contrary, the NR depicts a northward current in the eastern side of the channel and also in the west, where it is more intense. In the middle of the channel (0.62E-0.85E), there is a strip of weak northward current neither present in the CR. Top two panels of Fig. 4 shows the mean circulation in the region for the NR and CR.
@@
The data assimilation system employed is the Local Multimodel EnOI scheme previously described that was validated to correctly assimilate HFR observations in Hernandez-Lasheras et al. (2021). We here use the nudging initialization method after analysis, as it is the one employed in the WMOP operational system and it is less prone to produce discontinuities in the field which could affect the computation of FSLE, that will be later presented.
@@ 2.3 OSSE Set-up: Pseudo-Observations
For our experiment, the satellite and Argo pseudo observations have been extracted at the same position and time as the real observations in the previous chapter’s OSE (Hernandez-Lasheras et al., 2021), that is considered as reference. We simulated along-track sea level anomalies (SLA) from four different altimeters (Cryosat, Jason-2, Saral Altika, and HY-2). NR fields are interpolated in space and time to each satellite observation after removing the mean dynamic topography. For SST, we emulate the SST Foundation product, which does not account for the diurnal cycles, sub-sampling surface temperature fields from the NR at 8 a.m., with a 10 km resolution. The Argo profiles were sampled by interpolating the temperature and salinity fields in space and time. We added noise to every observation, which was randomly generated for each observation considering a Gaussian probability distribution with a standard deviation of the value of the error. The observation error has been considered the same as for the real observations. Table 2 indicates the value of the representation and instrumental errors considered for the different observations. For Argo observations only the horizontal representation error is shown. Note that the total error has the expression $\sigma^{2}_{tot}=\sigma^{2}_{rep}+\sigma^{2}_{ins}$.
@@
The histogram of the innovations (observation - model) is shown in Fig. 5 for the OSE (blue) and OSSE (red), whose results are also synthesized in Table 3. We can observe that all observations follow a similar distribution both for OSE and OSSE. The only discrepancy is found in the SLA, where we can observe a bias of 0.07 m in the OSE. This is a known issue concerning the value of the satellite SLA observations. Satellite SLA observations are computed by extracting the value of the absolute dynamic topography measured by the satellite to the mean dynamic topography (MDT). The MDT of the model generally differs from the MDT of the observations. Moreover, two different simulations have two different MDTs. The rise of the mean sea level generates a climatological trend in the observations which has not been corrected and thus, SLA observations from recent years present a positive bias which also affects the innovations. However, we believe this does not significantly affect the assimilation of SLA, as it is not corrected during the simulation (as discussed in the previous chapter (Hernandez-Lasheras et al., 2021)) and do not impact the geostrophic circulation. The values of the innovations standard deviation, which is directly related to the centered-root-mean-square-deviation (CRMSD), has small differences between OSE and OSSE for all observing sources.
@@
For the HFR observations we have followed a slightly different approach. We have considered two polygons, one containing the actual coverage area and another considering the potential future coverage that a set of two antennas settled in the western part of the IC might provide, according to expert criteria (Fig. 1). Within these areas, we have sub-sampled the daily mean velocity fields of the NR in a spatial grid of 3 km resolution, which corresponds to the HFR total (u-v) observing resolution in the area (Tintoré et al., 2012). We randomly dismissed 15% of the observations for simulating potential gaps in the antennas coverage. Again, we have introduced a Gaussian noise to the observations and considered the same instrumental and representativity error we previously used with the real data. Fig. 6 shows the histogram of the innovations for both variables of velocity during the two periods run.
@@ 3 Results
First, we assess the results of our experiments comparing them to the ones obtained in the previous chapter’s OSEs used as reference (Hernandez-Lasheras et al., 2021). When the forecast error reduction obtained with both perspectives (OSE and OSSE) is similar, we can consider the system is valid, and reliably estimates the impact of other potential observations. This would enhance our reliability in the assessment of the future observation system design. For this validation we analyze the period spanning from 21 September to 20 October from an Eulerian perspective, in a similar way as it was done in the previous OSE: a) For SLA, we compare for each day of simulation the model equivalents against the NR at every location of all along-track possible observations in the region; b) SST comparison between model and NR is performed at every observation point within a grid of 10 km resolution, like the one used to generate the pseudo-observations; c) Temperature and salinity fields are interpolated in time and space to the Argo float profile observations; d) For comparison against HFR data we interpolate the surface average fields to the position of the real observations. Note that in the OSSE perspective, we compare against the value of the true ocean state, represented by the NR simulation. Thus, although the evaluation with the NR is done at some of the same points from where we extract the assimilated observations, the validation can not be considered as fully independent.
@@
A further analysis of the results is shown in the Taylor diagrams in Fig. 7. For SLA, the model error decreases around 40% and the correlation between model and observation increases from 0.38 to 0.75 when using DA. Results are almost equal for all three simulations using DA. For SST, the comparison between model and NR is performed at every observation point within a grid of 10 km resolution, like the one used to generate the pseudo-observations. Results show how the error reduction is around 36%, while correlation increases for all simulations from 0.87 to 0.94. These results are of the order of magnitude of the ones obtained for real observations (shown in the previous OSE). Similarly, for the Argo T-S observations, the results obtained for the OSSE are very similar and of same the order of those obtained with real-world observations. The error is reduced by 41 and a 36 %, for temperature and salinity respectively, and the correlation increases in both cases (Fig. 8).
@@ 3.2 Impact of the HFR system expansion: Eulerian validation
We further assess the impact of OSSEs on the surface currents comparing the model fields in each simulation against the NR in the area of the IC. Note that the approach is different to what we previously did to validate the OSSE by comparing it to the OSE as only the data at the observing points during the coincident period were evaluated previously. The area used for the assessment in this section, which can be seen in Fig. 1, covers the entire IC, being wider than the coverage of the future HFR system. This way, we evaluate the impact of HFR DA beyond the coverage of the antennas and its effects on the transport in the region.
@@
During the second period (1-30 August), the meridional currents were less intense than for the other simulation period, both in NR and CR. Furthermore, in NR, the currents in the western part of the IC were northward, contrary to CR and the area’s mean dynamics. This in an anomalous situation circulation in the area, but that is known to happen few times every year. We aim to explore the potential impact of the HFR system extension in such situations, where the model could differ more from the observations. GNR degrades the forecasting of surface currents compared to CR. HFR observations are also needed to reduce the error and increase the correlation between the model simulations and the NR. When using HFR observations for the zonal velocity the correlation increases from 0.50 for the CR to 0.76 and 0.80, and the error is reduced by 23 and 29%, for H-A and H-F respectively.

### v2

@@ Abstract
The impact of the expansion of a high-frequency radar (HFR) system in a dynamic coastal area (the Ibiza Channel in the Western Mediterranean Sea) is evaluated through an Observing System Simulation Experiment (OSSE). The installation of two new antennas in the Iberian Peninsula would complement the existing ones in the islands of Ibiza and Formentera, providing surface currents observations of the full channel. Two different configurations of the same model, validated to give realistic simulations, are used: i) a Nature Run (NR) which is considered as the real ocean state and that is used to generate pseudo-observations, and ii) a Control Run (CR) in which the pseudo-observations are assimilated. The OSSE is first validated by comparison against a previous Observing System Experiment (OSE). The impact of the new antennas for forecasting surface currents is evaluated in two different periods with different levels of agreement between NR and CR. The HFR expansion is found to contribute to significantly correct the circulation patterns in the Channel, leading to surface merdional velocity error reductions up to 19%. The effects on the transport in the area are also analyzed from a Lagrangian perspective, showing that DA can help to better represent the Lagrangian Coherent Structures present in the NR and constrain the ocean dynamics.
@@ 1 Introduction
Numerical models provide a complete view of the three dimensional structure of the ocean. However, they are inevitably affected by errors from parametrization of non resolved physical processes, discretization issues, or the lack of accurate forcing. To improve reliability, numerical models for operational purposes should be fed with observations through data assimilation. Observing System Experiments (OSEs) assimilating data are numerical experiments that can be designed to evaluate the capability of specific observing systems and eventually correct model forecasts simulations. Similarly, the potential impact of observing systems has to be evaluated to help design these systems. Observing System Simulation Experiments (OSSEs) can be performed to help to optimally design an OOS or a future campaign (Kourafalou et al., 2015).
OSSEs were first developed for the atmospheric science community, and over time, specific design criteria have been developed to ensure the realism of the assessments performed, as defined in Atlas (1997). In ocean studies, multiple OSSEs had been done, however, most of them did not use a full-fledged DA system approach for the evaluation. Generally, Kalman filters, empirical orthogonal functions (EOF) based or different interpolation methods were used to map the observations and reconstruct the ocean state (Guinehut et al., 2004; Ballabrera-Poy et al., 2007; Sakov and Oke, 2008). Following the procedure established for atmospheric studies (Hoffman and Atlas, 2016), Halliwell et al. (2014) applied them for the first time in the ocean, and in the last years, several studies have been performed following the criteria exposed in that work, as we will do here. For instance, Gasparin et al. (2019) performed an evaluation of the influence of the future deep Argo float network, and Benkiran et al. (2021) assessed the impact of the assimilation of data from the future SWOT satellite mission in a global-high-resolution model. Aydoğdu et al. (2016, 2018) also used this same approach to perform a design of observing systems in the Marmara and the Adriatic seas respectively.
@@
The validation of these Lagrangian techniques is generally limited. LCS computed from model simulations can be compared to those calculated from geostrophic currents derived from altimetry products (Hernández-Carrasco and Orfila, 2018) which suffer limitations when approaching the coast (Vignudelli et al., 2019; Pascual et al., 2013). Also from HFR measured currents (Hernández-Carrasco et al., 2018), that are limited to cover small coastal areas. The validation of LCS can be performed by comparison with active tracers, as chlorophyll filaments (Lehahn et al., 2007; Hernández-Carrasco et al., 2018), which by their nature can not depict the full structures present in the ocean; SST fronts (d’Ovidio et al., 2004), which are inferred from satellite products that can be affected by clouds; or fish stock concentrations (Díaz-Barroso et al., 2022), tracked seabirds or marine predators (Kai et al., 2009), that are difficult to monitor. Here we will take profit of the full ocean state knowledge supposed in the OSSE approach. The use of a NR model for the validation of the LCS computed from the model simulations implies a step forward to address the question of how data assimilation can help models to correct the circulation, especially in coastal areas.
This study complements the work presented in Hernandez-Lasheras et al. (2021), where a series of OSEs were performed to evaluate the impact of HFR DA on the correction of surface currents in the Ibiza Channel (IC). In that work, it was shown that using HFR observations together with satellite observations (altimetry and sea surface temperature) and Argo temperature and salinity profiles increased the model’s capability to forecast surface currents. In this work we will use the same set-up, from the Western Mediterranean OPerational modelling system (WMOP).
@@ 2.1 Study Area and HFR system
As part of the Mediterranean HFR observing network ((Lorente et al., 2022; Reyes et al., 2022), SOCIB operates since 2012 two CODAR HFR antennas in the islands of Ibiza and Formentera, measuring surface currents in an area up to 80 km far off the coast (Tintoré et al., 2012). Here we evaluate the potential impact that a couple of new antennas potentially installed in the Iberian peninsula, in the eastern part of Cape la Nao (Fig. 1), would have on the WMOP modelling system. By installing these new antennas, the IC HFR system will expand its coverage area covering the entire channel. The area, shown in Fig. 1 (red dashed line), is considered the most likely coverage in terms of total velocities (u-v) that a couple of antennas installed in the western side of the Channel would provide together with the present system.
Figure 1: WMOP model domain and bathymetry in the Western Mediterranean Sea. HFR present (black dashed) and potentially future (red dashed) coverage area over which observations have been simulated. The blue rectangle represents the area selected for the validation.
@@ 2.2 OSSE set-up: Simulations
The OSSE perspective requires two model simulations. Here we used a fraternal twin OSSE system approach (Halliwell et al., 2014), in which two different configurations of the same model are employed: i) a Nature Run (NR), which is considered to represent the true ocean and that will be used for validation and to extract the pseudo-observations, and (ii) a Control Run (CR) model in which we will evaluate the impact of assimilating different observing sources. We use the WMOP modelling system (Juza et al., 2016; Mourre et al., 2018; Hernandez-Lasheras and Mourre, 2018), which is a regional configuration of ROMS (Shchepetkin and McWilliams, 2005) for the Western Mediterranean Sea coupled with a Local Multimodel Ensemble Optimal Interpolation (EnOI) data assimilation engine. It spans from Gibraltar Strait in the west to Corsica and Sardinia straits in the east, with a horizontal resolution around 2 km and 32 vertical terrain-following sigma levels (resulting in a vertical resolution between 1 and 2m at the surface). The NR and CR models share the same configuration, parametrization and atmospheric forcing, but differ in their initial state and boundary conditions. The CR is a free-run hindcast simulation developed and evaluated in Mourre et al. (2018) and Aguiar et al. (2019). It uses the Copernicus Marine Forecasting System for the Mediterranean Sea (CMEMS MED-MFC), with a 1/16 horizontal resolution (Simoncelli et al., 2017), as initial and lateral boundary conditions. This initial and boudary condition configuration is the same as that used in the companion OSE that we will use for validation (Hernandez-Lasheras et al., 2021). NR, in contrast, uses the Mercator Glorys reanalysis global product, with a 1/12 horizontal resolution (CMEMS GLOBAL_REANALYSIS_PHY_001_030) and has also been validated to give realistic simulations, comparing against observational data from satellites and Argo buoys. The atmospheric forcing, common for both simulations, is provided every 3 hours at $1/20^{\circ}$ resolution by the Spanish Meteorological Agency (AEMET) through the HIRLAM model (Undén et al., 2002) and the bathymetry is derived from a 1’ database (Smith and Sandwell, 1997).
@@
To further explore the capabilities of the new antenna under different possible circulation patterns, we selected another period during which the surface circulation from the two simulations present larger differences in the study area. This other period is August 2014. In the CR simulation, the dynamics during August are similar to the following September-October period, with northward currents in the east side of the channel and southward in the west, as can be seen in Fig. 2. On the contrary, the NR depicts a northward current in the eastern side of the channel and also in the west, where it is more intense. In the middle of the channel (0.62E-0.85E), there is a strip of weak northward current neither present in the CR. Top two panels of Fig. 4 shows the mean circulation in the region for the NR and CR.
@@
The data assimilation system employed is the Local Multimodel EnOI scheme previously successfully used to assimilate satellite observations, temperature and salinity profiles from CTDs and gliders as well as surface velocities from the Ibiza Channel HFR (Hernandez-Lasheras and Mourre, 2018; Hernandez-Lasheras et al., 2021). We here use the nudging initialization method after analysis, as it is the one employed in the WMOP operational system and it is less prone to produce discontinuities in the field which could affect the computation of FSLE, that will be later presented.
@@ 2.3 OSSE Set-up: Pseudo-Observations
For our experiment, the satellite and Argo pseudo observations have been extracted at the same position and time as the real observations analyzed in the OSE presented in (Hernandez-Lasheras et al., 2021). We simulated along-track sea level anomalies (SLA) from four different altimeters (Cryosat, Jason-2, Saral Altika, and HY-2). NR fields are interpolated in space and time to each satellite observation after removing the mean dynamic topography. For SST, we emulate the SST foundation product, which does not account for the diurnal cycles, sub-sampling surface temperature fields from the NR at 8 a.m., with a 10 km resolution. The Argo profiles were sampled by interpolating the temperature and salinity fields in space and time. We added noise to every observation, which was randomly generated for each observation considering a Gaussian probability distribution with a standard deviation of the value of the error. The observation error has been considered the same as for the real observations. Table 2 indicates the value of the representation and instrumental errors considered for the different observations. For Argo observations only the horizontal representation error is shown. Note that the total error has the expression $\sigma^{2}_{tot}=\sigma^{2}_{rep}+\sigma^{2}_{ins}$.
@@
The histogram of the innovations (observation - model) is shown in Fig. 5 for the OSE (blue) and OSSE (red), whose results are also synthesized in Table 3. We can observe that all observations follow a similar distribution both for OSE and OSSE. The only discrepancy is found in the SLA, where we can observe a bias of 0.07 m in the OSE. This is a common issue linked to the mismatch between the mean dynamic topography (MDT) of the model and that of the observations. Moreover, two different simulations using different initial conditions generally also show differences in the MDT. However, we believe this does not significantly affect the assimilation of SLA, as it is not corrected during the simulation (as discussed in the previous chapter (Hernandez-Lasheras et al., 2021)) and do not impact the geostrophic circulation. The values of the innovations standard deviation, which is directly related to the centered-root-mean-square-deviation (CRMSD), has small differences between OSE and OSSE for all observing sources.
@@
For the HFR observations we have followed a slightly different approach. We have considered two polygons, one containing the actual coverage area and another considering the potential future coverage that a set of two antennas settled in the western part of the IC might provide, according to expert criteria (Fig. 1). Within these areas, we have sub-sampled the daily mean velocity fields of the NR in a spatial grid of 3 km resolution, which corresponds to the HFR total (u-v) observing resolution in the area (Tintoré et al., 2012). We randomly discarded 15% of the observations for simulating potential gaps in the antennas coverage. Again, we have introduced a Gaussian noise to the observations and considered the same instrumental and representativity error we previously used with the real data. Fig. 6 shows the histogram of the innovations for both variables of velocity during the two periods run.
@@ 3 Results
First, we assess the results of our experiments comparing them to the ones obtained in the companion OSEs (Hernandez-Lasheras et al., 2021). The evaluation system will be considered as valid (i.e. it reliably estimates the impact of potential observations) when the forecast error reduction obtained in both cases (OSE and OSSE) is similar. For this validation we analyze the period spanning from 21 September to 20 October from an Eulerian perspective, in a similar way as it was done in the previous OSE: a) For SLA, we compare for each day of simulation the model equivalents against the NR at every location of all along-track possible observations in the region; b) SST comparison between model and NR is performed at every observation point within a grid of 10 km resolution, like the one used to generate the pseudo-observations; c) Temperature and salinity fields are interpolated in time and space to the Argo float profile observations; d) For comparison against HFR data we interpolate the surface average fields to the position of the real observations. Note that in the OSSE perspective, we compare against the value of the true ocean state, represented by the NR simulation.
@@
A further analysis of the results is shown in the Taylor diagrams in Fig. 7. For SLA, the model error decreases around 40% and the correlation between model and observation increases from 0.38 to 0.75 when using DA. Results are almost equal for all three simulations using DA. For SST, the comparison between model and NR is performed at every observation point within a grid of 10 km resolution, like the one used to generate the pseudo-observations. Results show how the error reduction is around 36%, while correlation increases for all simulations from 0.87 to 0.94. These results are of the same order of magnitude as the ones obtained for real observations (shown in the previous OSE). Similarly, for the Argo T-S observations, the results obtained for the OSSE are very similar and of the same the same order as those obtained with real-world observations. The error is reduced by 41 and a 36 %, for temperature and salinity respectively, and the correlation increases in both cases (Fig. 8).
@@ 3.2 Impact of the HFR system expansion: Eulerian validation
We further assess the impact of OSSEs on the surface currents comparing the model fields in each simulation against the NR in the area of the IC. The area used for the assessment in this section, which can be seen in Fig. 1, covers the entire IC, being wider than the coverage of the future HFR system. This way, we evaluate the impact of HFR DA beyond the coverage of the antennas and its effects on the transport in the region.
@@
During the second period (1-30 August), the meridional currents were less intense than for the other simulation period, both in NR and CR. Furthermore, in NR, the currents in the western part of the IC were northward, contrary to the CR and the area’s mean circulation pattern. This situation highlights the highly dynamical nature of the Ibiza Channel and provides an interesting study case complementary to the first study period. We aim to explore the potential impact of the HFR system extension in such situations, where the model could significantly differ from the observations. GNR degrades the forecasting of surface currents compared to CR. HFR observations are also needed to reduce the error and increase the correlation between the model simulations and the NR. When using HFR observations for the zonal velocity the correlation increases from 0.50 for the CR to 0.76 and 0.80, and the error is reduced by 23 and 29%, for H-A and H-F respectively.